OpenStudy (anonymous):

Suppose the tip of the minute hand of a clock is two inches from the center of the clock. Determine the distance traveled by the tip of the minute hand if 20 minutes have passed. How do you set this up? (from the trig section on linear and angular velocity)

OpenStudy (anonymous):



OpenStudy (anonymous):

Thanks! But what formula did u use? I just want to know so that I know how to do these kinds of problems.

OpenStudy (anonymous):

Pi* d is the circumference of a circle. For a circle with a 2" radius it is:\[\pi (2+2) \]In 20 minutes, the minute hand completes 20/60 of a circle. The distance that the tip of the minuted hand moves is\[\frac{20}{60}*\pi ( 2+2)=\frac{4 \pi }{3}=4.19 \text{ inches}\] By the way the clock diagram shows a 2 inch hour hand.

OpenStudy (anonymous):

My bad. minuted is a misspelling of the word, minute. If the problem deals with the hour hand then the answer above should be divided by 12. The hour hand only moves 1/12 of a circle per hour.

OpenStudy (anonymous):

Oh thank you so much! Great explanation, it was much simpler than I thought :)

OpenStudy (anonymous):

Sorry for the bad labeling, maybe  looks better..

OpenStudy (anonymous):

Glad I could be of help.

OpenStudy (anonymous):

Over looked the second diagram which is excellent.

OpenStudy (anonymous):

Haha, thanks. :)